# FNODE: Flow-Matching for data-driven simulation of constrained multibody systems

### v3

@@ Abstract
Data-driven modeling of constrained multibody dynamics remains challenged by (i) the training cost of Neural ODEs, which typically require backpropagation through an ODE solver, and (ii) error accumulation in rollout predictions. We introduce a Flow-Matching Neural ODE (FNODE) framework that learns the acceleration mapping directly from trajectory data by supervising accelerations rather than integrated states, turning training into a supervised regression problem and eliminating the ODE-adjoint/solver backpropagation bottleneck. Acceleration targets are obtained efficiently via numerical differentiation using a hybrid fast Fourier transform (FFT) and finite-difference (FD) scheme. Kinematic constraints are enforced through coordinate partitioning: FNODE learns accelerations only for the independent generalized coordinates, while the dependent coordinates are recovered by solving the position-level constraint equations. We evaluate FNODE on single and triple mass–spring–damper systems, a double pendulum, a slider–crank with and without friction, a vehicle model, and a cart–pole, and compare against MBD-NODE, LSTM, and fully connected baselines. Across these benchmarks, FNODE achieves improved prediction accuracy and training/runtime efficiency, while maintaining constraint satisfaction through the partitioning procedure. Our code and scripts are released as open source to support reproducibility and follow-on research.
Keywords Multibody dynamics $\cdot$ Neural ODE $\cdot$ Constrained dynamics $\cdot$ Scientific machine learning
@@ 2.2.1 Neural Ordinary Differential Equation (NODE)
$\frac{d\mathbf{z}(t)}{dt}=f(\mathbf{z}(t),t;\mathbf{\Theta})\;,$ (2)
where $f:\mathbb{R}^{n_{z}}\times\mathbb{R}^{+}\to\mathbb{R}^{n_{z}}$ is a neural network parameterized by $\mathbf{\Theta}$. The future state $\mathbf{z}(t)$ is obtained by solving the initial value problem
$\mathbf{z}(t)=\mathbf{z}(0)+\int_{0}^{t}f(\mathbf{z}(\tau),\tau;\mathbf{\Theta}),d\tau=\Phi(\mathbf{z}(0),f,t),$ (3)
@@ 2.2.2 Flow-Matching in NODE
$f(\mathbf{z}(t),t;\mathbf{\Theta}):\mathbb{R}^{2n_{z}}\times\mathbb{R}^{n_{\bm{\mu}}}\to\mathbb{R}^{n_{z}}\;.$ (8)
In inference phase, the forward integration process maintains the same mathematical structure as traditional NODEs with the definition from Eq. (3) and Eq. (4).
$\bm{Z}(t,\bm{\mu})=\bm{Z}(0,\bm{\mu})+\int_{0}^{t}f(\bm{Z}(\tau,\bm{\mu});\ \mathbf{\Theta})d\tau=\Phi(\bm{Z}(0,\bm{\mu}),f,t).$ (9)
Figure 1: Discretized forward pass of the FNODE framework for a general multibody dynamics (MBD) system. The solid arrows represent the data flow during both training and inference. The dashed arrows indicate latent inference operations. The dotted arrows denote the calculation of time derivative (training target). During training, “ground-truth” accelerations are computed from trajectory data using FFT- or finite-difference (FD)-based differentiation. Once trained, the neural model $f(\cdot;\bm{\Theta})$ predicts accelerations that are supplied to an ODE solver during inference to generate the system’s time evolution.
@@ 2.3.2 Finite Difference Method
For non-periodic datasets, finite-difference (FD) schemes provide a simple and robust alternative to FFT-based differentiation. Let $z^{\prime}_{t}$ denote the FD estimate of $z^{\prime}(t)$. Using the central-difference formula, $z^{\prime}_{t}=\tfrac{z(t+\Delta t)-z(t-\Delta t)}{2\Delta t}$, and thus $z^{\prime}(t)=z^{\prime}_{t}+\mathcal{O}(\Delta t^{2})$ [48]. At the boundaries, we use one-sided formulas, $z^{\prime}_{t}=\tfrac{z(t+\Delta t)-z(t)}{\Delta t}$ and $z^{\prime}_{t}=\tfrac{z(t)-z(t-\Delta t)}{\Delta t}$, and hence $z^{\prime}(t)=z^{\prime}_{t}+\mathcal{O}(\Delta t)$ in either case. Although less accurate than spectral differentiation, FD methods are more stable for short trajectories and at dataset boundaries, making them complementary in our hybrid FFT–FD approach.
Higher-order FD discretizations are also available. For example, let $z^{\prime(5)}_{t}$ denote the five-point central FD estimate,
$z^{\prime(5)}_{t}=\tfrac{-z(t+2\Delta t)+8z(t+\Delta t)-8z(t-\Delta t)+z(t-2\Delta t)}{12\Delta t}\;,$
so that $z^{\prime}(t)=z^{\prime(5)}_{t}+\mathcal{O}(\Delta t^{4})$, together with matching higher-order one-sided stencils near boundaries. In our setting, trajectories are short and derivative labels are obtained from sampled data; wider stencils increase the number of samples affected by boundary treatment and can amplify high-frequency noise. We therefore adopt the three-point central stencil in the interior and one-sided stencils at the endpoints as a conservative accuracy–robustness tradeoff.
To obtain a practical estimate of differentiation error on a given dataset, one may use two complementary consistency checks. First, compute derivative estimates with two stencil orders and define a relative disagreement indicator $\eta^{\prime}=\tfrac{\|z^{\prime(5)}-z^{\prime(3)}\|_{2}}{\|z^{\prime(5)}\|_{2}}$, where $z^{\prime(3)}$ and $z^{\prime(5)}$ denote the three-point and five-point FD estimates of $z^{\prime}(t)$ (and analogously for $z^{\prime\prime}$), and $\|\cdot\|_{2}$ denotes the discrete $\ell_{2}$ norm over the sampled trajectory. Second, define a reconstruction-consistency indicator by integrating the differentiated signals and comparing to the original trajectory, e.g., $\tilde{z}_{n+1}=\tilde{z}_{n}+\Delta t\,z^{\prime}_{n}$ with $\tilde{z}_{0}=z_{0}$, and form the normalized reconstruction residual $\eta_{z}=\tfrac{\|z-\tilde{z}\|_{2}}{\|z\|_{2}}$. These diagnostics provide a practical way to quantify sensitivity to the differentiation scheme and to detect differentiation artifacts when constructing acceleration targets for FNODE training; a systematic evaluation of $\eta^{\prime}$ and $\eta_{z}$ across datasets is deferred to future work and falls outside the scope of the present contribution.
@@ 2.4 Qualitative Analysis of Error Propagation in Integration
To connect integration error with training error and differentiation-induced label error, we decompose the one-step defect. Let the true augmented dynamics be $\dot{Z}=F(Z)$ and let $\widehat{F}_{\Theta}(Z)$ denote the learned vector field (FNODE). Let $\Psi_{\Delta t}(\cdot;F)$ denote one step of a numerical integrator with step size $\Delta t$ applied to the vector field $F$. In this view, $\Phi(Z)=\Psi_{\Delta t}(Z;F)$ and $\Phi_{\Theta}(Z)=\Psi_{\Delta t}(Z;\widehat{F}_{\Theta})$ represent the corresponding one-step maps. The true and learned rollouts satisfy $Z_{n+1}=\Psi_{\Delta t}(Z_{n};F)$ and $\widehat{Z}_{n+1}=\Psi_{\Delta t}(\widehat{Z}_{n};\widehat{F}_{\Theta})$, and the global error $E_{n}=\widehat{Z}_{n}-Z_{n}$ evolves as $E_{n+1}=\Psi_{\Delta t}(Z_{n}+E_{n};\widehat{F}_{\Theta})-\Psi_{\Delta t}(Z_{n};F)$.
@@
Adding and subtracting $\Psi_{\Delta t}(Z_{n};\widehat{F}_{\Theta})$ yields $E_{n+1}=\delta_{n}+\bigl(\Psi_{\Delta t}(Z_{n}+E_{n};\widehat{F}_{\Theta})-\Psi_{\Delta t}(Z_{n};\widehat{F}_{\Theta})\bigr)$, where $\delta_{n}:=\Psi_{\Delta t}(Z_{n};\widehat{F}_{\Theta})-\Psi_{\Delta t}(Z_{n};F)$ is the one-step defect at the true state. For an order-$p$ integrator, the defect admits a bound of the form $\|\delta_{n}\|\leq C_{\mathrm{int}}\Delta t^{p+1}+C_{\mathrm{fld}}\Delta t\,\|\widehat{F}_{\Theta}(Z_{n})-F(Z_{n})\|$, i.e., it combines the local truncation error of the time integrator and the effect of vector-field mismatch over one step.
@@
The learned-field error has two contributions. FNODE is trained on numerically differentiated acceleration labels $\tilde{a}=a+e_{\mathrm{diff}}$, so $\widehat{F}_{\Theta}-F=(\widehat{F}_{\Theta}-\tilde{F})+(\tilde{F}-F)$, where $\tilde{F}$ is the vector field consistent with the differentiated labels. The first term is controlled by the training objective; in particular, if $L(\Theta)=\tfrac{1}{N}\sum_{i}\|\widehat{F}_{\Theta}(Z_{i})-\tilde{F}(Z_{i})\|_{2}^{2}$, then $\varepsilon_{\mathrm{train}}:=\sqrt{L(\Theta)}$ provides an RMS proxy for the learned-field mismatch on the training distribution. The second term reflects differentiation artifacts and can be assessed through the diagnostics introduced in Section 2.3.2, such as the stencil-disagreement indicator $\eta^{\prime}$ and the trajectory reconstruction residual $\eta_{z}$.
@@
Finally, the second term in $E_{n+1}$ captures error propagation through integration. If the integrator is stable and satisfies $\bigl\|\tfrac{\partial}{\partial Z}\Psi_{\Delta t}(Z;\widehat{F}_{\Theta})\bigr\|\leq 1+C\Delta t$ over the rollout region, then $\|\Psi_{\Delta t}(Z_{n}+E_{n};\widehat{F}_{\Theta})-\Psi_{\Delta t}(Z_{n};\widehat{F}_{\Theta})\|\leq(1+C\Delta t)\|E_{n}\|$, which implies the recursion $\|E_{n+1}\|\leq(1+C\Delta t)\|E_{n}\|+\|\delta_{n}\|$. Consequently, $\|E_{n}\|$ grows at most like $(1+C\Delta t)^{n}$, and for $T=n\Delta t$ one obtains $\|E(T)\|\lesssim\exp(CT)\bigl(\|E(0)\|+\sum_{k}\|\delta_{k}\|\bigr)$. This shows explicitly how integration error accumulation couples to (i) the integrator order and step size through $C_{\mathrm{int}}\Delta t^{p+1}$, (ii) the training error through $\varepsilon_{\mathrm{train}}$, and (iii) differentiation-induced label error through the diagnostics $\eta^{\prime}$ and $\eta_{z}$.

### v4

@@ Abstract
Data-driven modeling of constrained multibody dynamics remains challenged by (i) the training cost of Neural ODEs, which typically require backpropagation through an ODE solver, and (ii) error accumulation in rollout predictions. We introduce a Flow-Matching Neural ODE (FNODE) framework that learns the acceleration mapping directly from trajectory data by supervising accelerations rather than integrated states, turning training into a supervised regression problem and eliminating the ODE-adjoint/solver backpropagation bottleneck. Acceleration targets are obtained efficiently via numerical differentiation using a hybrid fast Fourier transform (FFT) and finite-difference (FD) scheme. Kinematic constraints are enforced through coordinate partitioning: FNODE learns accelerations only for the independent generalized coordinates, while the dependent coordinates are recovered by solving the position-level constraint equations. We evaluate FNODE on single and triple mass–spring–damper systems, a double pendulum, a slider–crank with and without friction, a vehicle model, and a cart–pole, and compare against MBD-NODE, LSTM, and fully connected baselines. We also compared the error contributions between numerical differentiation and Neural Network learned field mismatch in single-mass-spring-damper system, which proves the error of learned acceleration field is dominated by the error of Neural Network and confirms the robustness of numerical differentiation. Across these benchmarks, FNODE achieves improved prediction accuracy and training/runtime efficiency, while maintaining constraint satisfaction through the partitioning procedure. Our code and scripts are released as open source to support reproducibility and follow-on research.
Keywords Multibody dynamics $\cdot$ Neural ODE $\cdot$ Constrained dynamics $\cdot$ Friction dynamics $\cdot$ Scientific machine learning $\cdot$ Flow-Matching
MSC2020: 68T07, 68T45
@@ 2.2.1 Neural Ordinary Differential Equation (NODE)
$\frac{d\mathbf{z}(t)}{dt}=f_{\Theta}(\mathbf{z}(t),t)\;,$ (2)
where $f_{\Theta}:\mathbb{R}^{n_{z}}\times\mathbb{R}^{+}\to\mathbb{R}^{n_{z}}$ is a neural network parameterized by $\mathbf{\Theta}$. The future state $\mathbf{z}(t)$ is obtained by solving the initial value problem
$\mathbf{z}(t)=\mathbf{z}(0)+\int_{0}^{t}f_{\Theta}(\mathbf{z}(\tau),\tau),d\tau=\Phi(\mathbf{z}(0),f_{\Theta},t),$ (3)
@@ 2.2.2 Flow-Matching in NODE
$f_{\Theta}(\mathbf{z}(t),\bm{\mu}):\mathbb{R}^{2n_{z}}\times\mathbb{R}^{n_{\bm{\mu}}}\to\mathbb{R}^{n_{z}}\;.$ (8)
In the inference phase, the forward integration process maintains the same mathematical structure as traditional NODEs with the definition from Eq. (3) and Eq. (4).
$\bm{Z}(t,\bm{\mu})=\bm{Z}(0,\bm{\mu})+\int_{0}^{t}f_{\Theta}(\bm{Z}(\tau,\bm{\mu}))d\tau=\Phi(\bm{Z}(0,\bm{\mu}),t)\circ f_{\Theta}=\Phi_{\Theta}(\bm{Z}(0,\bm{\mu}),t).$ (9)
Figure 1: Discretized forward pass of the FNODE framework for a general multibody dynamics (MBD) system. The solid arrows represent the data flow during both training and inference. The dashed arrows indicate latent inference operations. The dotted arrows denote the calculation of time derivative (training target). During training, “ground-truth” accelerations are computed from trajectory data using FFT- or finite-difference (FD)-based differentiation. Once trained, the neural model $f_{\Theta}(\cdot)$ predicts accelerations that are supplied to an ODE solver during inference to generate the system’s time evolution.
@@ 2.3.2 Finite Difference Method
For non-periodic datasets, finite-difference (FD) schemes provide a simple and robust alternative to FFT-based differentiation. Let $\dot{z}_{t}$ denote the FD estimate of $\dot{z}(t)$. Using the central-difference formula, $\dot{z}_{t}=\tfrac{z(t+\Delta t)-z(t-\Delta t)}{2\Delta t}$, and thus $\dot{z}(t)=\dot{z}_{t}+\mathcal{O}(\Delta t^{2})$ [18]. At the boundaries, we use one-sided formulas, $\dot{z}_{t}=\tfrac{z(t+\Delta t)-z(t)}{\Delta t}$ and $\dot{z}_{t}=\tfrac{z(t)-z(t-\Delta t)}{\Delta t}$, and hence $\dot{z}(t)=\dot{z}_{t}+\mathcal{O}(\Delta t)$ in either case. Although less accurate than spectral differentiation, FD methods are more stable for short trajectories and at dataset boundaries, making them complementary in our hybrid FFT–FD approach.
Higher-order FD discretizations are also available. For example, let $\dot{z}^{(5)}_{t}$ denote the five-point central FD estimate,
$\dot{z}^{(5)}_{t}=\tfrac{-z(t+2\Delta t)+8z(t+\Delta t)-8z(t-\Delta t)+z(t-2\Delta t)}{12\Delta t}\;,$
so that $\dot{z}(t)=\dot{z}_{t}^{(5)}+\mathcal{O}(\Delta t^{4})$, together with matching higher-order one-sided stencils near boundaries. In our setting, trajectories are short and derivative labels are obtained from sampled data; wider stencils increase the number of samples affected by boundary treatment and can amplify high-frequency noise. We therefore adopt the three-point central stencil in the interior and one-sided stencils at the endpoints as a conservative accuracy–robustness tradeoff.
To obtain a practical estimate of differentiation error on a given dataset, one may use two complementary consistency checks. First, compute derivative estimates with two stencil orders and define a relative disagreement indicator $\eta^{\prime}=\tfrac{\|\dot{z}_{t}^{(5)}-\dot{z}_{t}^{(3)}\|_{2}}{\|\dot{z}_{t}^{(5)}\|_{2}}$, where $\dot{z}_{t}^{(3)}$ and $\dot{z}_{t}^{(5)}$ denote the three-point and five-point FD estimates of $\dot{z}(t)$ (and analogously for $\ddot{z}(t)$), and $\|\cdot\|_{2}$ denotes the discrete $\ell_{2}$ norm over the sampled trajectory. Second, define a reconstruction-consistency indicator by integrating the differentiated signals and comparing to the original trajectory, e.g., $\tilde{z}_{t+\Delta t}=\tilde{z}_{t}+\Delta t\,\dot{z}_{t}$ with $\tilde{z}_{0}=z_{0}$, and form the normalized reconstruction residual $\eta_{z}=\tfrac{\|z_{t}-\tilde{z}_{t}\|_{2}}{\|z_{t}\|_{2}}$. These diagnostics provide a practical way to quantify sensitivity to the differentiation scheme and to detect differentiation artifacts when constructing acceleration targets for FNODE training; a systematic evaluation of $\eta^{\prime}$ and $\eta_{z}$ across datasets is deferred to future work and falls outside the scope of the present contribution.
@@ 2.4 Qualitative Analysis of Error Propagation in Integration
To connect integration error with training error and differentiation-induced label error, we decompose the one-step defect. Let the true augmented dynamics be $\ddot{z}=f(Z)$ and let $\hat{f}_{\Theta}(Z)$ denote the learned vector field (FNODE trained on numerically differentiated acceleration labels). Let $\Psi_{\Delta t}(\cdot;f)$ denote one step of a numerical integrator with step size $\Delta t$ applied to the vector field $f$. In this view, $\Phi(Z)=\Psi_{\Delta t}(Z;f)$ and $\Phi_{\Theta}(Z)=\Psi_{\Delta t}(Z;\hat{f}_{\Theta})$ represent the corresponding one-step maps. The true and learned rollouts satisfy $Z_{n+1}=\Psi_{\Delta t}(Z_{n};f)$ and $\widehat{Z}_{n+1}=\Psi_{\Delta t}(\widehat{Z}_{n};\hat{f}_{\Theta})$, and the global error $E_{n}=\widehat{Z}_{n}-Z_{n}$ evolves as $E_{n+1}=\Psi_{\Delta t}(Z_{n}+E_{n};\hat{f}_{\Theta})-\Psi_{\Delta t}(Z_{n};f)$.
@@
Adding and subtracting $\Psi_{\Delta t}(Z_{n};\hat{f}_{\Theta})$ yields $E_{n+1}=\delta_{n}+\bigl(\Psi_{\Delta t}(Z_{n}+E_{n};\hat{f}_{\Theta})-\Psi_{\Delta t}(Z_{n};\hat{f}_{\Theta})\bigr)$, where $\delta_{n}:=\Psi_{\Delta t}(Z_{n};\hat{f}_{\Theta})-\Psi_{\Delta t}(Z_{n};f)$ is the one-step defect at the true state. Under standard smoothness assumptions on the true dynamics and assuming the use of a consistent order-$p$ numerical integrator whose induced numerical flow satisfies a Lipschitz-type stability bound over the rollout region, the one-step defect can be bounded as: $\|\delta_{n}\|\leq C_{\mathrm{int}}\Delta t^{p+1}+C_{\mathrm{fld}}\Delta t\,\|\hat{f}_{\Theta}(Z_{n})-f(Z_{n})\|$, i.e., it combines the local truncation error of the time integrator and the effect of vector-field mismatch over one step.
@@
The learned-field error has two contributions. FNODE is trained on numerically differentiated acceleration labels $\tilde{\ddot{z}}=\ddot{z}+e_{\mathrm{diff}}$, so $\hat{f}_{\Theta}-f=(\hat{f}_{\Theta}-\tilde{f})+(\tilde{f}-f)$, where $\tilde{f}$ is the vector field consistent with the differentiated labels. The first term is controlled by the training objective; in particular, if $L(\Theta)=\tfrac{1}{N}\sum_{i}\|\hat{f}_{\Theta}(Z_{i})-\tilde{f}(Z_{i})\|_{2}^{2}$, then $\varepsilon_{\mathrm{train}}:=\sqrt{L(\Theta)}$ provides an RMS proxy for the learned-field mismatch on the training distribution. The second term reflects differentiation artifacts and can be assessed through the diagnostics introduced in Section 2.3.2, such as the stencil-disagreement indicator $\eta^{\prime}$ and the trajectory reconstruction residual $\eta_{z}$.
@@
Finally, the second term in $E_{n+1}$ captures error propagation through integration. If the integrator is stable and satisfies $\bigl\|\tfrac{\partial}{\partial Z}\Psi_{\Delta t}(Z;\hat{f}_{\Theta})\bigr\|\leq 1+C\Delta t$ over the rollout region, then $\|\Psi_{\Delta t}(Z_{n}+E_{n};\hat{f}_{\Theta})-\Psi_{\Delta t}(Z_{n};\hat{f}_{\Theta})\|\leq(1+C\Delta t)\|E_{n}\|$, which implies the recursion $\|E_{n+1}\|\leq(1+C\Delta t)\|E_{n}\|+\|\delta_{n}\|$. Consequently, $\|E_{n}\|$ grows at most like $(1+C\Delta t)^{n}$, and for $T=n\Delta t$ one obtains $\|E(T)\|\lesssim\exp(CT)\bigl(\|E(0)\|+\sum_{k}\|\delta_{k}\|\bigr)$. This shows explicitly how integration error accumulation couples to (i) the integrator order and step size through $C_{\mathrm{int}}\Delta t^{p+1}$, (ii) the training error through $\varepsilon_{\mathrm{train}}$, and (iii) differentiation-induced label error through the diagnostics $\eta^{\prime}$ and $\eta_{z}$. For the three error sources, (i) could be illustrated by the trajectory error, while (ii) and (iii) represent the vector-field mismatch in acceleration. These three error sources are further analyzed below.
2.4.1 Integrator error
For a $p$th-order accurate time integrator, the global trajectory-level MSE is represented as:
$\widetilde{\mathrm{MSE}}_{Z}(\Delta t)=\frac{1}{N}\sum_{n=0}^{N-1}\big\|\Phi(Z_{n})-Z_{n+1}\big\|^{2},$
where $\Phi(Z_{n})$ is the one-step update produced using the ground truth acceleration with step size $\Delta t$. As mentioned above, the single-step truncation error is $\mathcal{O}(\Delta t^{p+1})$, so the error over $t\in[0,T]$ is $\mathcal{O}(\Delta t^{p})$ uniformly, then
$\mathrm{MSE}_{Z}(\Delta t)\sim\mathcal{O}(\Delta t^{2p}),$
Note that since RK4 has $p=4$, the global discretization error (in the state) over a fixed time horizon typically scales as $\mathcal{O}(\Delta t^{4})$ under standard smoothness assumptions.
2.4.2 Neural Network learned field mismatch
We train FNODE with analytical accelerations $\ddot{z}$ and acceleration calculated by FD $\tilde{\ddot{z}}$ as targets and evaluate the predicted accelerations $\mathring{\ddot{z}}_{n}$ (trained on $\ddot{z}$) and $\hat{\ddot{z}}$ (trained on $\tilde{\ddot{z}}$) on the ground truth $\ddot{z}_{n}$, reporting $\mathring{\mathrm{MSE}}_{\ddot{z}}(\Delta t)$ and $\widehat{\mathrm{MSE}}_{\ddot{z}}(\Delta t)$:
$\mathring{\mathrm{MSE}}_{\ddot{z}}(\Delta t)=\frac{1}{N}\sum_{n=0}^{N-1}\|\mathring{\ddot{z}}_{n}-\ddot{z}_{n}\|^{2}$
$\widehat{\mathrm{MSE}}_{\ddot{z}}(\Delta t)=\frac{1}{N}\sum_{n=0}^{N-1}\|\hat{\ddot{z}}_{n}-\ddot{z}_{n}\|^{2}$
2.4.3 Differentiation-induced label error
To interpret FD slopes, we partition indices into an interior set $\mathcal{I}$ and a boundary set $\mathcal{B}=\{0,N-1\}$, where $|\mathcal{B}|=2$ is a constant independent of $N$ (the samples where one-sided stencils are used) and $\mathcal{I}=\{1,\dots,N-2\}$, where $|\mathcal{I}|=N-2$. Define the pointwise FD acceleration error $\tilde{e}_{n}=\tilde{\ddot{z}}_{n}-\ddot{z}_{n}$. Let $q_{\mathcal{I}}$ and $q_{\mathcal{B}}$ denote the order of the finite-difference method used to compute accelerations in the interior and boundary regions, respectively, i.e., $\|\tilde{e}_{n}\|=\mathcal{O}(\Delta t^{q_{\mathcal{I}}})$ for $n\in\mathcal{I}$ and $\|\tilde{e}_{n}\|=\mathcal{O}(\Delta t^{q_{\mathcal{B}}})$ for $n\in\mathcal{B}$. Here $\mathcal{I}$ contains $N-\mathcal{O}(1)$ interior time indices, whereas $\mathcal{B}$ contains only $\mathcal{O}(1)$ boundary indices (e.g., the first/last few samples required by the stencil). Over a fixed time horizon $T$, the number of time samples satisfies $N\sim T/\Delta t$.
Then the interior and boundary contributions to acceleration MSE scale as
$\widetilde{\mathrm{MSE}}_{\ddot{z},\mathcal{I}}(\Delta t)$ $=\frac{1}{N}\sum_{n\in\mathcal{I}}\|\tilde{e}_{n}\|^{2}$ (24)
$\sim\frac{|\mathcal{I}|}{N}\,\mathcal{O}_{\mathcal{I}}(\Delta t^{2q_{\mathcal{I}}})=\frac{N-2}{N}\mathcal{O}_{\mathcal{I}}(\Delta t^{2q_{\mathcal{I}}})=(1-\frac{2}{N})\mathcal{O}_{\mathcal{I}}(\Delta t^{2q_{\mathcal{I}}}),$
$\widetilde{\mathrm{MSE}}_{\ddot{z},\mathcal{B}}(\Delta t)=\frac{1}{N}\sum_{n\in\mathcal{B}}\|\tilde{e}_{n}\|^{2}\sim\frac{|\mathcal{B}|}{N}\,\mathcal{O}_{\mathcal{B}}(\Delta t^{2q_{\mathcal{B}}})=\frac{2}{N}\mathcal{O}_{\mathcal{B}}(\Delta t^{2q_{\mathcal{B}}})$ (25)
Since $N-1=\frac{T}{\Delta t}$ and the physical time is not extremely long ($10^{6}$ or above), $\frac{1}{N-1}=\frac{\Delta t}{T}\sim\mathcal{O}(\Delta t)$. Besides, as $N\gg 1$,
$\frac{2}{N}\approx\frac{2}{N-1}=2\frac{\Delta t}{T}\sim\mathcal{O}(\Delta t)$
Then we can rewrite Eq.(25) as:
$\widetilde{\mathrm{MSE}}_{\ddot{z},\mathcal{B}}(\Delta t)=\frac{1}{N}\sum_{n\in\mathcal{B}}\|\tilde{e}_{n}\|^{2}\sim\frac{|\mathcal{B}|}{N}\,\mathcal{O}_{\mathcal{B}}(\Delta t^{2q_{\mathcal{B}}})=\frac{2}{N}\mathcal{O}_{\mathcal{B}}(\Delta t^{2q_{\mathcal{B}}})\approx\mathcal{O}_{\mathcal{B}}(\Delta t^{2q_{\mathcal{B}}+1})$ (26)
The global FD acceleration MSE is:
$\widetilde{\mathrm{MSE}}_{\ddot{z}}(\Delta t)=\widetilde{\mathrm{MSE}}_{\ddot{z},\mathcal{I}}(\Delta t)+\widetilde{\mathrm{MSE}}_{\ddot{z},\mathcal{B}}(\Delta t),$ (27)
The overall MSE in Eq. (27) is dominated, for sufficiently small $\Delta t$, by the term with the smaller exponent in $\Delta t$ (i.e., the term that decays more slowly and therefore yields a larger error). Therefore, according to Eq. (24) and Eq. (26), if $2q_{\mathcal{I}}<2q_{\mathcal{B}}+1$, the interior error dominates, and vice versa.
To simplify step-size convergence order comparisons, the root mean squared error (RMSE) can also be used to quantify the error. Although RMSE makes it difficult to directly separate the contributions from $\mathcal{I}$ and $\mathcal{B}$, its effect on the observed convergence order is universal. Therefore, the order of $\widetilde{\mathrm{RMSE}}_{\ddot{z},\mathcal{I}}(\Delta t)$ is $q_{\mathcal{I}}$, whereas the order of $\widetilde{\mathrm{RMSE}}_{\ddot{z},\mathcal{B}}(\Delta t)$ is $q_{\mathcal{B}}+\tfrac{1}{2}$.
For a fixed-order time integrator with a fixed step size, any acceleration error, regardless of its source, enters the same numerical update operator at each step. Therefore, when comparing rollouts generated with the same integrator and $\Delta t$, differences observed at the trajectory level primarily reflect differences in the underlying acceleration (vector-field) errors, rather than differences in the integrator. Consequently, to assess the robustness of our numerical differentiation, it is necessary to compare, under identical integration settings, the learned vector-field mismatch against the numerical-differentiation error.
The quantitative error-propagation analysis that isolates the contribution of each source is provided in Section 3.2.